OpenStudy (anonymous):

can someone solve: tan^-1(tan(pi/4-x/2))=? -pi/2<=x<=3pi/2

OpenStudy (anonymous):

shld it be pi/4-x/2

OpenStudy (jhannybean):

\[\large \tan^{-1}(\tan(x)) = x\] This rule applies to cosines and sine functions as well. The inverse function always outdoes the function itself.

OpenStudy (anonymous):

so what do you say my answer is right or not

OpenStudy (jhannybean):

You are absolutely right.

OpenStudy (anonymous):

then u know range of inverse tan (-pi/2,pi/2)

OpenStudy (anonymous):

now x/2 has range [-pi/4,3pi/4] as per my question

OpenStudy (jhannybean):

range of inverse tangent = \(\large (\frac{\pi}{2} , \frac{-\pi}{2})\)

OpenStudy (anonymous):

ohk i got myself cleared i was confused between tan^-1(x) and tan^-1(tanx) as for both range will varry thanks a lot Jhannybean

OpenStudy (jhannybean):

Good job, :)